OpenStudy (anonymous):

Diana received 65 points on a project for school. She can make changes and receive three-tenths of the missing points back. She can make as many corrections as she wants. Create the formula for the sum of this geometric series and explain your steps in solving for the maximum grade Diana can receive. Identify this as converging or diverging. okay so I know it's convergent and i'm going to assume there are 35 missing points and a1 = 65 and the rate is 3/10 but i'm not sure how to set up the equation?

OpenStudy (anonymous):

This is simplistic. Use u-substitution to find the initial state of a qubit before it is observed in a plus/minus basis, then use Grover's Algorithm to search through a lattice of N-indexes.

OpenStudy (anonymous):

After that you should end up with a 78 variable diffy-q.

OpenStudy (anonymous):

Resolve and convert notation:

OpenStudy (anonymous):

i haven't learned about grover's algorithm or any of that stuff yet so i don't understand what you mean

OpenStudy (anonymous):

Well, you're teacher must suck. This is obviously a reference to differential calc

OpenStudy (anonymous):

Lel I'm just breaking your balls. @SyedaLovesPie

OpenStudy (anonymous):

lol i'm in algebra 2 right now so i haven't learned about calc yet @happytales

OpenStudy (anonymous):

....

OpenStudy (anonymous):

Ohhh excuses excuses :P... I taught myself calc in middle school, starting with multivariable. Pertaining to your question, aren't geometric series just ratios expressed as a constant?

OpenStudy (anonymous):

ahhh yeah i'm not that good at math haha. and i don't know what you mean? basically i'm just trying to find the total amount of points she can receive after correcting all of the problems she missed

OpenStudy (anonymous):

wait do you think if i do 65 - 35(.3) / 1 - .3 i'll get the right answer?

OpenStudy (anonymous):

Hmm... it wants you to express an equation using sigma notation I assume. So assume the first term is \[Sn=65\] So assume the equation: \[\sum_{k=0}^{4-1}(Sr^k)=a (\frac{ 1-r^n }{ 1-r })\] Following?

OpenStudy (anonymous):

for the most part but where did you get 4-1 from?

OpenStudy (anonymous):

ffs n-1. Ignore that.

OpenStudy (anonymous):

So your expressed equation: \[\sum_{k=0}^{35-1}(65*0.3)=65(\frac{ 1-0.3^35 }{ 1-0.3) }\] Ignore the random 5... the carat didn't want to work -_- should be 35.

OpenStudy (anonymous):

ohhh okay thank you so much!